# BCWildfire: A Long-term Multi-factor Dataset and Deep Learning Benchmark for Boreal Wildfire Risk Prediction

### v1

@@ Abstract
Code and Datasetshttps://github.com/SynUW/mmFire
@@ Introduction
Dataset Lookback Forecast Temporal
Length Horizon Resolution
FireCube Daily Daily Daily
WildfireDB Daily Daily Daily
WildfireSpreadTS 14-94 days 4-94 days Daily
Next Day Daily Daily Daily
Wildfire Spread
SeasFire Cube / / 8 days
mmFire (ours) $\leq$ 25 years $\leq$ 25 years Daily
Table 1: Comparison between current next day or event-based wildfire prediction datasets and mmFire.
@@
To address these fundamental limitations, we present mmFire, a comprehensive spatiotemporal dataset that covers British Columbia, Canada, and adjacent regions, spanning approximately 240 million hectares over the period 2000 to 2024 and incorporating 38 multimodal wildfire driving factors. We further conduct a systematic benchmark of deep learning architectures, including CNN-based, linear-based, Transformer-based, and Mamba-based models, for the next day wildfire risk prediction to evaluate their effectiveness on this challenging task.
@@
  • We introduce mmFire, a large-scale spatiotemporal dataset for wildfire risk prediction, consisting of 2.4 million samples across 240 million hectares of boreal landscapes over a 25-year period. The dataset integrates 38 multimodal wildfire-driving factors, addressing major limitations in existing resources related to geographic coverage, feature completeness, and long-term temporal representation.
  • In contrast to existing datasets that primarily focus on wildfire spread and behavior modeling based on spatial information, mmFire is specifically designed to support time series forecasting. This design enables the modeling of cumulative interactions among wildfire drivers over time, thereby facilitating the development of models with enhanced temporal reasoning capabilities and the potential to establish new benchmarks for wildfire prediction.
@@
Overall, the number and spatiotemporal coverage of publicly available wildfire datasets remain limited. Existing datasets tend to focus on fire spread modeling for individual events within short time windows, primarily in the U.S. and Mediterranean regions. This limited scope constrains the development of wildfire ignition risk prediction models, which rely on long-term accumulative drivers and are crucial for proactive wildfire management (XU2025632).. Moreover, due to the high sensitivity of wildfire behavior to local environmental and climatic conditions, models trained on region-specific or structurally constrained datasets often fail to generalize to boreal forests (gerard2023wildfirespreadts).
@@
Figure 2: Dataset overview: including fuel, fire detection (red box), topography (yellow box), human activities (black box), and meteorological factors (blue box).
@@ Methods
mmFire Benchmark Dataset
@@
Overall, we construct a long-term time series forecasting dataset with consistent spatiotemporal resolution, covering British Columbia, Canada, and surrounding regions from 2000 to 2024. Each sample includes 38 daily covariates grouped into five categories: vegetation (fuel) conditions, active fire detections, meteorological factors, human activity, and topographic features, as shown in Figure 2. The dataset primarily sources data from MODIS satellite remote sensing products, ERA5-Land reanalysis data (munoz2021era5), OpenStreetMap, and the ASTER digital elevation model (DEM) (jarvis2004practical). As illustrated in Figure 2, each day’s observation is organized into a structured data cube with all features aligned. Given the varying native resolutions of the data sources, we resample all inputs to a uniform spatial resolution of 1 km and a daily temporal resolution to ensure consistency across both dimensions.
@@
Topographical factors also play an important role in the ignition and spread of wildfires (coen2013wrf). Steep slopes help fire to rise quickly uphill, increasing the rate of spread; aspect affects sunlight exposure and dryness, thereby influencing the flammability of fuels; and hillshade reflects, to some extent, the shading of terrain on solar radiation, indirectly affecting local microclimate and burning conditions (deng2025daily). Therefore, we selected 30-meter resolution ASTER DEM data and calculated slope, aspect, and hillshade based on it (ASTERGDEMv3_2019). Additionally, water bodies often serve as natural barriers that can significantly reduce the probability of fire spread, so we also calculated the distance to the nearest water body using OpenStreetMap’s water distribution map, further enhancing the model’s ability to represent fire behavior (OpenStreetMap2025).
@@
Human activities represent a major source of wildfire ignition (bowring2024road). To capture their potential influence, we incorporated the yearly 500 m resolution MODIS Land Use product (MCD12Q1) to account for the impact of different land use types on wildfire occurrence, as substantial contrasts exist between agricultural and forested areas in terms of fuel characteristics and human intervention (friedl2019mcd12q1). In addition, using settlement and power grid distribution data from OpenStreetMap, we computed the distance to the nearest infrastructure for each pixel. Since human activities tend to cluster around infrastructure, such proximity may increase the likelihood of ignition events (OpenStreetMap2025).
@@
Following the aforementioned data collection and processing procedures, our final dataset for model training, validation, and testing comprises 1,015,275 samples, including 338,425 positive samples and 676,850 negative samples. To examine the distributional characteristics and temporal dynamics of wildfire-driving factors across multiple time scales, we computed the probability density functions for wildfire and non-wildfire pixels over the preceding 365 days and analyzed their temporal trends during the last 10 days prior to ignition, as shown in Figure 3(a) and Figure 3(b), respectively. The results indicate that temperature increases, fuel accumulation, precipitation decreases, and surface soil moisture reductions are strongly associated with wildfire ignition. However, substantial distributional overlap remains between burned and unburned areas, underscoring the intrinsic stochasticity of wildfire occurrence in space and time—under similar environmental conditions, fires may either ignite or remain dormant within high-risk zones. Interestingly, regions experiencing wildfire ignition tend to exhibit slightly higher deep-layer (100–289 cm) soil water content, suggesting that wildfires often occur in ecologically moist, fuel-rich environments where deep soil moisture does not directly mitigate surface drying and ignition risk.
@@
We evaluated ten time series forecasting models on the BCWildfire dataset, covering a diverse range of architectures, including CNN-based, linear-based, Transformer-based, and Mamba-based approaches. Specifically, the models include the CNN-based SCINet (liu2022scinet), linear-based models such as CrossLinear (zhou2025crosslinear), and TSMixer (wang2024timexer), Transformer-based models including Corssformer (zhang2023crossformer) and FEDFormer (zhou2022fedformer), as well as the Mamba-based S_Mamba (wang2025mamba).
@@
Performance Ceilings.
Our benchmark reveals persistent performance ceilings among existing wildfire prediction architectures under the 10 day input configuration. Even the best performing models, including CrossLinear (recall 87.56%), Crossformer (87.49%), and FEDformer (91.19%), plateau below 92% recall, while precision remains within a narrow 83–88% range. These results underscore fundamental challenges in wildfire modeling, such as extreme class imbalance, strong spatiotemporal heterogeneity, and complex nonlinear interactions among drivers.
@@ Conclusion
We present mmFire, a 25 year multimodal wildfire benchmark dataset covering 240 million hectares across British Columbia and adjacent northern regions. The dataset integrates 38 key wildfire driving factors, including meteorological, fuel, topographic, and human activity variables, harmonized at 1 km spatial and daily temporal resolution. Based on this resource, we establish the first unified time series forecasting benchmark for wildfire risk prediction.

### v2

@@ Abstract
Code and Datasetshttps://github.com/SynUW/BCWildfire
@@ Introduction
Dataset Temporal Span (Lookback/Forecast) Resolution
FireCube Daily/Daily Daily
WildfireDB Daily/Daily Daily
WildfireSpreadTS 14–94 d/4–94 d Daily
Next Day Wildfire Spread Daily/Daily Daily
SeasFire Cube N/A/N/A 8 d
BCWildfire (ours) $\leq$$\leq25 y$ Daily
Table 1: Comparison between current next day or event-based wildfire prediction datasets and BCWildfire.
@@
To address these fundamental limitations, we present BCWildfire, a comprehensive spatiotemporal dataset that covers British Columbia, Canada, and adjacent regions, spanning approximately 240 million hectares over the period 2000 to 2024 and incorporating 38 multimodal wildfire driving factors. We further conduct a systematic benchmark of deep learning architectures, including CNN-based, linear-based, Transformer-based, and Mamba-based models, for the next day wildfire risk prediction to evaluate their effectiveness on this challenging task.
@@
  • We introduce BCWildfire, a large-scale spatiotemporal dataset for wildfire risk prediction, consisting of 2.4 million samples across 240 million hectares of boreal landscapes over a 25-year period. The dataset integrates 38 multimodal wildfire-driving factors, addressing major limitations in existing resources related to geographic coverage, feature completeness, and long-term temporal representation.
  • In contrast to existing datasets that primarily focus on wildfire spread and behavior modeling based on spatial information, BCWildfire is specifically designed to support time series forecasting. This design enables the modeling of cumulative interactions among wildfire drivers over time, thereby facilitating the development of models with enhanced temporal reasoning capabilities and the potential to establish new benchmarks for wildfire prediction.
@@
Overall, the number and spatiotemporal coverage of publicly available wildfire datasets remain limited. Existing datasets tend to focus on fire spread modeling for individual events within short time windows, primarily in the U.S. and Mediterranean regions. This limited scope constrains the development of wildfire ignition risk prediction models, which rely on long-term accumulative drivers and are crucial for proactive wildfire management (XU2025632). Moreover, due to the high sensitivity of wildfire behavior to local environmental and climatic conditions, models trained on region-specific or structurally constrained datasets often fail to generalize to boreal forests (gerard2023wildfirespreadts).
@@
Figure 2: Partial overview of the dataset, illustrating fuel (green), fire detection (red), topography (yellow), human activities (black), and meteorological factors (blue).
@@ Methods
BCWildfire Benchmark Dataset
@@
Overall, we construct a long-term time series forecasting dataset with consistent spatiotemporal resolution, covering British Columbia, Canada, and surrounding regions from 2000 to 2024. Each sample includes 38 daily covariates grouped into five categories: vegetation (fuel) conditions, active fire detections, meteorological factors, human activity, and topographic features, as shown in Figure 2. The dataset primarily sources data from MODIS satellite remote sensing products, ERA5-Land reanalysis data (munoz2021era5), OpenStreetMap (openstreetmap2021), and the ASTER digital elevation model (DEM) (jarvis2004practical). As illustrated in Figure 2, each day’s observation is organized into a structured data cube with all features aligned. Given the varying native resolutions of the data sources, we resample all inputs to a uniform spatial resolution of 1 km and a daily temporal resolution to ensure consistency across both dimensions.
@@
Topographical factors also play an important role in the ignition and spread of wildfires (coen2013wrf). Steep slopes help fire to rise quickly uphill, increasing the rate of spread; aspect affects sunlight exposure and dryness, thereby influencing the flammability of fuels; and hillshade reflects, to some extent, the shading of terrain on solar radiation, indirectly affecting local microclimate and burning conditions (deng2025daily). Therefore, we selected 30-meter resolution ASTER DEM data and calculated slope, aspect, and hillshade based on it (jarvis2004practical). Additionally, water bodies often serve as natural barriers that can significantly reduce the probability of fire spread, so we also calculated the distance to the nearest water body using OpenStreetMap’s water distribution map, further enhancing the model’s ability to represent fire behavior (openstreetmap2021).
@@
Human activities represent a major source of wildfire ignition (bowring2024road). To capture their potential influence, we incorporated the yearly 500 m resolution MODIS Land Use product (MCD12Q1) to account for the impact of different land use types on wildfire occurrence, as substantial contrasts exist between agricultural and forested areas in terms of fuel characteristics and human intervention (friedl2019mcd12q1). In addition, using settlement and power grid distribution data from OpenStreetMap, we computed the distance to the nearest infrastructure for each pixel. Since human activities tend to cluster around infrastructure, such proximity may increase the likelihood of ignition events (openstreetmap2021).
@@
Following the aforementioned data collection and processing procedures, our final dataset for model training, validation, and testing comprises 1,015,275 samples, including 338,425 positive samples and 676,850 negative samples. To examine the distributional characteristics and temporal dynamics of wildfire-driving factors across multiple time scales, we computed the probability density functions for wildfire and non-wildfire pixels over the preceding 365 days and analyzed their temporal trends during the last 10 days prior to ignition, as shown in Figure 3(a) and Figure 3(b), respectively. The results indicate that temperature increases, fuel accumulation, precipitation decreases, and surface soil moisture reductions are strongly associated with wildfire ignition. However, substantial distributional overlap remains between burned and unburned areas, underscoring the intrinsic stochasticity of wildfire occurrence in space and time, whereby under similar environmental conditions, fires may either ignite or remain dormant even within high-risk zones. Interestingly, regions experiencing wildfire ignition tend to exhibit slightly higher deep-layer (100–289 cm) soil water content, suggesting that wildfires often occur in ecologically moist, fuel-rich environments where deep soil moisture does not directly mitigate surface drying and ignition risk.
@@
We evaluated 6 time series forecasting models on the BCWildfire dataset, covering a diverse range of architectures, including CNN-based, linear-based, Transformer-based, and Mamba-based approaches. Specifically, the models include the CNN-based SCINet (liu2022scinet), linear-based models such as CrossLinear (zhou2025crosslinear), and TSMixer (wang2024timexer), Transformer-based models including Corssformer (zhang2023crossformer) and FEDFormer (zhou2022fedformer), as well as the Mamba-based S_Mamba (wang2025mamba).
@@
Performance Ceilings
Our benchmark reveals persistent performance ceilings among existing wildfire prediction architectures under the 10 day input configuration. Even the best performing models, including CrossLinear (recall 87.56%), Crossformer (87.49%), and FEDformer (91.19%), plateau below 92% recall, while precision remains within a narrow 83–88% range. These results underscore fundamental challenges in wildfire risk prediction, such as extreme class imbalance, strong spatiotemporal heterogeneity, and complex nonlinear interactions among drivers.
@@ Conclusion
We present BCWildfire, a 25 year multimodal wildfire benchmark dataset covering 240 million hectares across British Columbia and adjacent northern regions. The dataset integrates 38 key wildfire driving factors, including meteorological, fuel, topographic, and human activity variables, harmonized at 1 km spatial and daily temporal resolution. Based on this resource, we establish the first unified time series forecasting benchmark for wildfire risk prediction.